f(×)=6/ײ3×-10




a. X=-2



b. x= 0



c. x=5

1
Expert's answer
2022-05-30T15:10:38-0400
"f(x)=\dfrac{6}{x^2-3x-10}"

"x^2-3x-10\not=0"

"(x+2)(x-5)\not=0"

"x\not=-2, x\not=5"

Domain: "(-\infin, -2)\cap (-2, 5)\cap(5, \infin)"

a.


"f(-2)=\dfrac{6}{(-2)^2-3(-2)-10} , undefined"

"f(-2)" is undefined.


b.


"f(0)=\dfrac{6}{(0)^2-3(0)-10}=-\dfrac{3}{5}"

"f(0)=-\dfrac{3}{5}" .


c.


"f(5)=\dfrac{6}{(5)^2-3(5)-10} , undefined"

"f(5)" is undefined.